LISTSERV mailing list manager LISTSERV 16.0

Help for HEANET-FAULTS Archives


HEANET-FAULTS Archives

HEANET-FAULTS Archives


February 2008


View:

Show Author | Hide Author
Show Table of Contents | Hide Table of Contents

 

Table of Contents:

HEA-NOC/20070904-1 [UPDATE] modsecurity install on samhain.heanet.ie (1 message)
HEA-NOC/20071022-4 [CLOSE] Installation of new DC rectifier in CWT PoP (1 message)
HEA-NOC/20071219-5 [CLOSE] ATM link to Limerick flapped (1 message)
HEA-NOC/20071231-3 [REOPEN] v6 Routing Issues to a & b.gtld-servers.net (1 message)
HEA-NOC/20071231-3 [UPDATE] v6 Routing Issues to a & b.gtld-servers.net (1 message)
HEA-NOC/20080107-6 [CLOSE] HEAnet Network Freeze (30th Jan ~ 4th Feb 2008) (1 message)
HEA-NOC/20080112-1 [CLOSE] Multiple flaps on connecting circuit to BFCE (1 message)
HEA-NOC/20080121-12 [CLOSE] ESBT Planned Outage Notice (1 message)
HEA-NOC/20080121-9 [CLOSE] IOS Upgrades (1 message)
HEA-NOC/20080121-9 [UPDATE] IOS Upgrades (1 message)
HEA-NOC/20080124-4 [CLOSE] Janet connection down (1 message)
HEA-NOC/20080124-4 [UPDATE] Janet connection down (3 messages)
HEA-NOC/20080126-5 [UPDATE] Bluenet Dublin ring brief loss of resilience (1 message)
HEA-NOC/20080127-5 [CLOSE] St. Angela's brief flap (1 message)
HEA-NOC/20080129-10 [CLOSE] Upgrade to SNMP module on centralised DC Rectifier (1 message)
HEA-NOC/20080129-10 [UPDATE] Upgrade to SNMP module on centralised DC Rectifier (3 messages)
HEA-NOC/20080129-9 [UPDATE] WIT migration to new IP equipment (2 messages)
HEA-NOC/20080131-6 [CLOSE] Transition UCC To New IP Network (1 message)
HEA-NOC/20080131-6 [UPDATE] Transition UCC To New IP Network (2 messages)
HEA-NOC/20080201-3 [CLOSE] Multiple flaps on backbone link from Limerick to Cork (1 message)
HEA-NOC/20080201-3 [OPEN] Multiple flaps on backbone link from Limerick to Cork (1 message)
HEA-NOC/20080201-3 [UPDATE] Multiple flaps on backbone link from Limerick to Cork (2 messages)
HEA-NOC/20080201-6 [CLOSE] Scheduled outage on STACS link (1 message)
HEA-NOC/20080201-6 [OPEN] Scheduled outage on STACS link (1 message)
HEA-NOC/20080202-1 [OPEN] Commodity IP service maintenance (1 message)
HEA-NOC/20080202-3 [OPEN] Hermes connectivity loss (1 message)
HEA-NOC/20080202-3 [UPDATE] Hermes connectivity loss (3 messages)
HEA-NOC/20080204-2 [CLOSE] St. Angela's Link Instability & Loss of Connectivity (1 message)
HEA-NOC/20080204-2 [OPEN] St. Angela's LInk Instability & Loss of Connectivity (1 message)
HEA-NOC/20080204-2 [UPDATE] St. Angela's Link Instability & Loss of Connectivity (2 messages)
HEA-NOC/20080204-5 [CLOSE] IADT Migration to New IP Equipment (1 message)
HEA-NOC/20080204-5 [OPEN] IADT Migration to New IP Equipment (1 message)
HEA-NOC/20080204-5 [UPDATE] IADT Migration to New IP Equipment (1 message)
HEA-NOC/20080204-8 [CLOSE] Loss of Connectivity to IADT & Reslience to NUIM & GEANT (1 message)
HEA-NOC/20080204-8 [OPEN] Loss Of Reslience to NUIM (1 message)
HEA-NOC/20080204-8 [UPDATE] Loss of Connectivity to IADT & Loss Of Reslience to NUIM (1 message)
HEA-NOC/20080204-8 [UPDATE] Loss of Connectivity to IADT & Reslience to NUIM & GEANT (1 message)
HEA-NOC/20080205-1 [CLOSE] IPv6 root server addresses (1 message)
HEA-NOC/20080205-1 [OPEN] IPv6 root server addresses (1 message)
HEA-NOC/20080205-1 [UPDATE] IPv6 root server addresses (1 message)
HEA-NOC/20080205-13 [OPEN] Geant Primary Link Down (1 message)
HEA-NOC/20080205-13 [UPDATE] Geant Primary Link Down (4 messages)
HEA-NOC/20080205-18 [CLOSE] Connectivity to Carlow College lost (1 message)
HEA-NOC/20080205-18 [OPEN] Connectivity to Carlow College lost (1 message)
HEA-NOC/20080205-18 [UPDATE] Connectivity to Carlow College lost (1 message)
HEA-NOC/20080205-2 [CLOSE] Temporary Loss of Routing to Citywest Core Router (1 message)
HEA-NOC/20080205-2 [OPEN] Temporary Loss of Routing to Citywest Core Router (1 message)
HEA-NOC/20080205-2 [UPDATE] Temporary Loss of Routing to Citywest Core Router (2 messages)
HEA-NOC/20080205-8 [CLOSE] Package updates and kernel upgrades on multiple machines (1 message)
HEA-NOC/20080205-8 [OPEN] Package updates and kernel upgrades on multiple machines (1 message)
HEA-NOC/20080205-8 [UPDATE] Package updates and kernel upgrades on multiple machines (5 messages)
HEA-NOC/20080206-10 [OPEN] Multiple flaps on HRB link (1 message)
HEA-NOC/20080206-10 [UPDATE] Multiple flaps on HRB link (1 message)
HEA-NOC/20080206-14 [OPEN/CLOSE] HEAnet mail server problem (1 message)
HEA-NOC/20080206-3 [CLOSE] Loss of Connection to IADT (1 message)
HEA-NOC/20080206-3 [OPEN] Loss of Connection to IADT (1 message)
HEA-NOC/20080206-3 [UPDATE] Loss of Connection to IADT (1 message)
HEA-NOC/20080206-4 [OPEN] Enabling Netflow on CRS (1 message)
HEA-NOC/20080206-4 [UPDATE] Enabling Netflow on CRS (4 messages)
HEA-NOC/20080207-1 [CLOSE] Loss of Connection to St Angela's (1 message)
HEA-NOC/20080207-1 [OPEN] Loss of Connection to St Angela's (1 message)
HEA-NOC/20080207-1 [UPDATE] Loss of Connection to St Angela's (2 messages)
HEA-NOC/20080207-11 [CLOSE] Loss of signal on adva equipment between Kilcarbery and Citywest (1 message)
HEA-NOC/20080207-11 [OPEN] Loss of signal on adva equipment between Kilcarbery and Citywest (1 message)
HEA-NOC/20080207-12 [OPEN] GÉANT replacement STM-16 (1 message)
HEA-NOC/20080207-12 [UPDATE] GÉANT replacement STM-16 (6 messages)
HEA-NOC/20080207-2 [CLOSE] Power Failure in TCD PoP (1 message)
HEA-NOC/20080207-2 [OPEN] Power Failure in TCD PoP (1 message)
HEA-NOC/20080208-11 [CLOSE] Loss of Geant Secondary BGP Peering (1 message)
HEA-NOC/20080208-11 [OPEN] Loss of Geant Secondary BGP Peering (1 message)
HEA-NOC/20080208-13 [CLOSE] Unable to divert NOC number to answering service (1 message)
HEA-NOC/20080208-13 [OPEN] Unable to divert NOC number to answering service (1 message)
HEA-NOC/20080208-2 [CLOSE] Transition CIT To New IP Network (1 message)
HEA-NOC/20080208-2 [OPEN] Transition CIT To New IP Network (1 message)
HEA-NOC/20080208-2 [UPDATE] Transition CIT To New IP Network (2 messages)
HEA-NOC/20080208-4 [CLOSE] Cessation of legacy ITnet Commodity IP service (1 message)
HEA-NOC/20080208-4 [OPEN] Cessation of legacy ITnet Commodity IP service (1 message)
HEA-NOC/20080208-4 [UPDATE] Cessation of legacy ITnet Commodity IP service (1 message)
HEA-NOC/20080209-1 [CLOSE] Connectivity issues to UL (1 message)
HEA-NOC/20080209-1 [UPDATE] Connectivity issues to UL (1 message)
HEA-NOC/20080209-3 [CLOSE] NUI, Primary Link Down (1 message)
HEA-NOC/20080209-3 [OPEN] NUI, Primary Link Down (1 message)
HEA-NOC/20080209-3 [UPDATE] NUI, Primary Link Down (3 messages)
HEA-NOC/20080209-5 [OPEN/CLOSE] Loss of connectivity to DIAS (1 message)
HEA-NOC/20080209-6 [CLOSE] Connectivity issues to IUA (1 message)
HEA-NOC/20080209-6 [OPEN] Connectivity issues to IUA (1 message)
HEA-NOC/20080209-6 [UPDATE] Connectivity issues to IUA (4 messages)
HEA-NOC/20080209-7 [CLOSE] Connectivity issues to BCFE (1 message)
HEA-NOC/20080209-7 [OPEN] Connectivity issues to BCFE (1 message)
HEA-NOC/20080210-2 [OPEN] No connectivity to cdveckillester1 (1 message)
HEA-NOC/20080210-2 [UPDATE] No connectivity to cdveckillester1 (6 messages)
HEA-NOC/20080211-6 [CLOSE] Installation of 10G cards on all INEX switchs (1 message)
HEA-NOC/20080211-6 [OPEN] Installation of 10G cards on all INEX switchs (1 message)
HEA-NOC/20080212-11 [OPEN] Report of slow connectivity from Cork School of Music (1 message)
HEA-NOC/20080212-11 [UPDATE] Report of slow connectivity from Cork School of Music (1 message)
HEA-NOC/20080212-2 [CLOSE] Test listserv failover (1 message)
HEA-NOC/20080212-2 [OPEN] Test listserv failover (1 message)
HEA-NOC/20080212-2 [UPDATE] Test listserv failover (2 messages)
HEA-NOC/20080212-4 [CLOSE] Transition LIT To New IP Network (1 message)
HEA-NOC/20080212-4 [OPEN] Transition LIT To New IP Network (1 message)
HEA-NOC/20080212-4 [UPDATE] Transition LIT To New IP Network (2 messages)
HEA-NOC/20080212-6 [CLOSE] Loss of connectivity to GDA (1 message)
HEA-NOC/20080212-6 [OPEN] Loss of connectivity to GDA (1 message)
HEA-NOC/20080213-1 [CLOSE] The servers hosting the video streaming service are down (1 message)
HEA-NOC/20080213-1 [OPEN] The servers hosting the video streaming service are down (1 message)
HEA-NOC/20080213-1 [UPDATE] The servers hosting the video streaming service are down (1 message)
HEA-NOC/20080213-3 [OPEN] ESBT Plan to carry out repairs in the Kilbarry/Killonan POPs (1 message)
HEA-NOC/20080213-4 [CLOSE] Global Crossing Maintenance (1 message)
HEA-NOC/20080213-4 [OPEN] Outage to replace router hardware (1 message)
HEA-NOC/20080213-6 [CLOSE] IT Carlow Commodity IP Service Maintenance (1 message)
HEA-NOC/20080213-6 [OPEN] IT Carlow Commodity IP Service Maintenance (1 message)
HEA-NOC/20080214-9 [OPEN] ESB Planned Maintenance affecting Kilbarry/Killoteran Fibre (1 message)
HEA-NOC/20080215-1 [OPEN] Fault repair on fibres in the Claremorris-Ballina duct (1 message)
HEA-NOC/20080215-3 [OPEN] ESBT Plan to carry out repairs in the Kilbarry/Killoteran POPs (1 message)
HEA-NOC/20080215-6 [CLOSE] Listserv and SixXS6 server upgrade (1 message)
HEA-NOC/20080215-6 [OPEN] Listserv and SixXS6 server upgrade (1 message)
HEA-NOC/20080215-6 [UPDATE] Listserv and SixXS6 server upgrade (2 messages)
HEA-NOC/20080215-8 [OPEN] Scheduled Power Outage at CIT on 22nd Feb (1 message)
HEA-NOC/20080218-8 [CLOSE] Planned power outage in WIT Carriganore (1 message)
HEA-NOC/20080218-8 [OPEN] Planned power outage in WIT Carriganore (1 message)
HEA-NOC/20080218-8 [UPDATE] Planned power outage in WIT Carriganore (2 messages)
HEA-NOC/20080219-2 [CLOSE] Circuits down to VEC Killester and Liberties (1 message)
HEA-NOC/20080219-2 [OPEN] Circuits down to VEC Killester and Liberties (1 message)
HEA-NOC/20080219-2 [UPDATE] Circuits down to VEC Killester and Liberties (2 messages)
HEA-NOC/20080219-3 [CLOSE] Hardware failure of samhain, web services server (1 message)
HEA-NOC/20080219-3 [OPEN] Hardware failure of samhain, web services server (1 message)
HEA-NOC/20080219-3 [UPDATE] Hardware failure of samhain, web services server (2 messages)
HEA-NOC/20080221-1 [OPEN] ATM Interface to St. Angela's flapping (1 message)
HEA-NOC/20080221-1 [UPDATE] ATM Interface to St. Angela's flapping (9 messages)
HEA-NOC/20080221-1 [UPDATE] STACS flapping circuit (5 messages)
HEA-NOC/20080221-16 [OPEN] BT fibre repair works at Manuella Junction (1 message)
HEA-NOC/20080221-8 [CLOSE] Configuration change on NCAD CPE to accommodate new network. (1 message)
HEA-NOC/20080221-8 [OPEN] Configuration change on NCAD CPE to accommodate new network. (1 message)
HEA-NOC/20080221-8 [UPDATE] Configuration change on NCAD CPE to accommodate new network. (2 messages)
HEA-NOC/20080223-1 [CLOSE] The network access swich in the UL Outreach centre rebooted (1 message)
HEA-NOC/20080223-1 [OPEN] The network access swich in the UL Outreach centre rebooted (1 message)
HEA-NOC/20080223-2 [CLOSE] A circuit to Mary Immaculate college transitioned briefly (1 message)
HEA-NOC/20080223-2 [OPEN] A circuit to Mary Immaculate college transitioned briefly (1 message)
HEA-NOC/20080224-1 [CLOSE] Leased lines to two clients in the Mount Street area are down (1 message)
HEA-NOC/20080224-1 [OPEN] Leased lines to two clients in the Mount Street area are down (1 message)
HEA-NOC/20080224-1 [UPDATE] Leased lines to two clients in the Mount Street area are down (1 message)
HEA-NOC/20080225-1 [CLOSE] Setup mysql replication between primary and standby webserver (1 message)
HEA-NOC/20080225-1 [OPEN] Setup mysql replication between primary and standby webserver (1 message)
HEA-NOC/20080225-10 [OPEN] IT Carlow Migration to New IP Equipment (1 message)
HEA-NOC/20080225-10 [UPDATE] IT Carlow Migration to New IP Equipment (1 message)
HEA-NOC/20080225-12 [CLOSE] Start up listserv on second virtual machine (1 message)
HEA-NOC/20080225-12 [OPEN] Start up listserv on second virtual machine (1 message)
HEA-NOC/20080225-8 [CLOSE] Commodity IP service maintenance (1 message)
HEA-NOC/20080225-8 [OPEN] Commodity IP service maintenance (1 message)
HEA-NOC/20080225-9 [CLOSE] AIT Migration to New IP Equipment (1 message)
HEA-NOC/20080225-9 [OPEN] AIT Migration to New IP Equipment (1 message)
HEA-NOC/20080225-9 [UPDATE] AIT Migration to New IP Equipment (1 message)
HEA-NOC/20080227-3 [OPEN] BT Maintenance (1 message)
HEA-NOC/20080227-4 [OPEN] BT Maintenance (1 message)
HEA-NOC/20080227-6 [OPEN] GTI (Georgia Tech Institute) Migration to New IP equipment (1 message)
HEA-NOC/20080227-8 [CLOSE] Scheduled power work in TCD PoP (1 message)
HEA-NOC/20080227-8 [OPEN] Scheduled power work in TCD PoP (1 message)
HEA-NOC/20080228-13 [CLOSE] Loss of Connectivity - HRB (1 message)
HEA-NOC/20080228-13 [OPEN] Loss of Connectivity - HRB (1 message)
HEA-NOC/20080228-4 [CLOSE] HEAnet Office Phone System Maintenance (1 message)
HEA-NOC/20080228-4 [OPEN] HEAnet Office Phone System Maintenance (1 message)
HEA-NOC/20080229-10 [OPEN] St Angela's link flapped (1 message)
HEA-NOC/20080229-4 [OPEN] Listserv services unavailable (1 message)
HEA-NOC/20080229-4 [UPDATE] Listserv services unavailable (2 messages)
HEA-NOC/20080229-9 [OPEN] Cessation of ITnet STM-1 (Cork) (1 message)
HEAnet Extended NOC Cover announcement... (1 message)
[OPEN] The servers hosting the video streaming service are down (1 message)
[UPDATE] Hardware failure of samhain, web services server (1 message)

LISTSERV Archives

LISTSERV Archives

HEANET-FAULTS Home

HEANET-FAULTS Home

Subject Sorted by Subject, Chronologically

From

Date

Size

New Thread

HEA-NOC/20070904-1 [UPDATE] modsecurity install on samhain.heanet.ie

HEA-NOC/20070904-1 [UPDATE] modsecurity install on samhain.heanet.ie

Aidan Carty <[log in to unmask]>

Wed, 27 Feb 2008 10:58:47 +0000

58 lines

New Thread

HEA-NOC/20071022-4 [CLOSE] Installation of new DC rectifier in CWT PoP

HEA-NOC/20071022-4 [CLOSE] Installation of new DC rectifier in CWT PoP

Gareth Eason <[log in to unmask]>

Thu, 7 Feb 2008 12:16:36 +0000

55 lines

New Thread

HEA-NOC/20071219-5 [CLOSE] ATM link to Limerick flapped

HEA-NOC/20071219-5 [CLOSE] ATM link to Limerick flapped

Gareth Eason <[log in to unmask]>

Thu, 21 Feb 2008 15:48:56 +0000

64 lines

New Thread

HEA-NOC/20071231-3 [REOPEN] v6 Routing Issues to a & b.gtld-servers.net

HEA-NOC/20071231-3 [REOPEN] v6 Routing Issues to a & b.gtld-servers.net

Brian Nisbet <[log in to unmask]>

Fri, 8 Feb 2008 09:52:29 +0000

47 lines

New Thread

HEA-NOC/20071231-3 [UPDATE] v6 Routing Issues to a & b.gtld-servers.net

HEA-NOC/20071231-3 [UPDATE] v6 Routing Issues to a & b.gtld-servers.net

Paul Cosgrove <[log in to unmask]>

Tue, 12 Feb 2008 12:10:59 +0000

51 lines

New Thread

HEA-NOC/20080107-6 [CLOSE] HEAnet Network Freeze (30th Jan ~ 4th Feb 2008)

HEA-NOC/20080107-6 [CLOSE] HEAnet Network Freeze (30th Jan ~ 4th Feb 2008)

Gareth Eason <[log in to unmask]>

Mon, 4 Feb 2008 17:42:28 +0000

57 lines

New Thread

HEA-NOC/20080112-1 [CLOSE] Multiple flaps on connecting circuit to BFCE

HEA-NOC/20080112-1 [CLOSE] Multiple flaps on connecting circuit to BFCE

Gareth Eason <[log in to unmask]>

Wed, 20 Feb 2008 12:34:38 +0000

61 lines

New Thread

HEA-NOC/20080121-12 [CLOSE] ESBT Planned Outage Notice

HEA-NOC/20080121-12 [CLOSE] ESBT Planned Outage Notice

Brian Nisbet <[log in to unmask]>

Wed, 6 Feb 2008 13:09:57 +0000

36 lines

New Thread

HEA-NOC/20080121-9 [CLOSE] IOS Upgrades

HEA-NOC/20080121-9 [CLOSE] IOS Upgrades

Garwin Liu <[log in to unmask]>

Tue, 12 Feb 2008 14:38:41 +0000

42 lines

New Thread

HEA-NOC/20080121-9 [UPDATE] IOS Upgrades

HEA-NOC/20080121-9 [UPDATE] IOS Upgrades

Paul Cosgrove <[log in to unmask]>

Wed, 6 Feb 2008 09:15:43 +0000

40 lines

New Thread

HEA-NOC/20080124-4 [CLOSE] Janet connection down

HEA-NOC/20080124-4 [CLOSE] Janet connection down

Brian Nisbet <[log in to unmask]>

Mon, 11 Feb 2008 09:16:35 +0000

57 lines

New Thread

HEA-NOC/20080124-4 [UPDATE] Janet connection down

HEA-NOC/20080124-4 [UPDATE] Janet connection down

Brian Nisbet <[log in to unmask]>

Fri, 8 Feb 2008 12:47:23 +0000

52 lines

HEA-NOC/20080124-4 [UPDATE] Janet connection down

Brian Nisbet <[log in to unmask]>

Mon, 4 Feb 2008 11:21:45 +0000

49 lines

HEA-NOC/20080124-4 [UPDATE] Janet connection down

Daniel Lete <[log in to unmask]>

Sat, 2 Feb 2008 08:10:10 +0000

45 lines

New Thread

HEA-NOC/20080126-5 [UPDATE] Bluenet Dublin ring brief loss of resilience

HEA-NOC/20080126-5 [UPDATE] Bluenet Dublin ring brief loss of resilience

Brian Boyle <[log in to unmask]>

Mon, 11 Feb 2008 15:54:34 +0000

34 lines

New Thread

HEA-NOC/20080127-5 [CLOSE] St. Angela's brief flap

HEA-NOC/20080127-5 [CLOSE] St. Angela's brief flap

Louis Twomey <[log in to unmask]>

Fri, 1 Feb 2008 16:17:08 +0000

56 lines

New Thread

HEA-NOC/20080129-10 [CLOSE] Upgrade to SNMP module on centralised DC Rectifier

HEA-NOC/20080129-10 [CLOSE] Upgrade to SNMP module on centralised DC Rectifier

Gareth Eason <[log in to unmask]>

Thu, 7 Feb 2008 12:02:17 +0000

52 lines

New Thread

HEA-NOC/20080129-10 [UPDATE] Upgrade to SNMP module on centralised DC Rectifier

HEA-NOC/20080129-10 [UPDATE] Upgrade to SNMP module on centralised DC Rectifier

Gareth Eason <[log in to unmask]>

Wed, 6 Feb 2008 10:03:11 +0000

49 lines

HEA-NOC/20080129-10 [UPDATE] Upgrade to SNMP module on centralised DC Rectifier

Gareth Eason <[log in to unmask]>

Wed, 6 Feb 2008 09:06:50 +0000

45 lines

HEA-NOC/20080129-10 [UPDATE] Upgrade to SNMP module on centralised DC Rectifier

Garreth Malone <[log in to unmask]>

Mon, 4 Feb 2008 18:10:25 +0000

43 lines

New Thread

HEA-NOC/20080129-9 [UPDATE] WIT migration to new IP equipment

HEA-NOC/20080129-9 [UPDATE] WIT migration to new IP equipment

Cormac O'Cianain <[log in to unmask]>

Wed, 13 Feb 2008 10:02:15 +0000

50 lines

HEA-NOC/20080129-9 [UPDATE] WIT migration to new IP equipment

Cormac O'Cianain <[log in to unmask]>

Thu, 7 Feb 2008 10:22:08 +0000

48 lines

New Thread

HEA-NOC/20080131-6 [CLOSE] Transition UCC To New IP Network

HEA-NOC/20080131-6 [CLOSE] Transition UCC To New IP Network

Brian Nisbet <[log in to unmask]>

Thu, 7 Feb 2008 10:57:31 +0000

50 lines

New Thread

HEA-NOC/20080131-6 [UPDATE] Transition UCC To New IP Network

HEA-NOC/20080131-6 [UPDATE] Transition UCC To New IP Network

Brian Nisbet <[log in to unmask]>

Wed, 6 Feb 2008 10:17:28 +0000

46 lines

HEA-NOC/20080131-6 [UPDATE] Transition UCC To New IP Network

Brian Nisbet <[log in to unmask]>

Fri, 1 Feb 2008 16:15:29 +0000

41 lines

New Thread

HEA-NOC/20080201-3 [CLOSE] Multiple flaps on backbone link from Limerick to Cork

HEA-NOC/20080201-3 [CLOSE] Multiple flaps on backbone link from Limerick to Cork

Louis Twomey <[log in to unmask]>

Fri, 1 Feb 2008 17:22:22 +0000

51 lines

New Thread

HEA-NOC/20080201-3 [OPEN] Multiple flaps on backbone link from Limerick to Cork

HEA-NOC/20080201-3 [OPEN] Multiple flaps on backbone link from Limerick to Cork

Louis Twomey <[log in to unmask]>

Fri, 1 Feb 2008 12:24:26 +0000

35 lines

New Thread

HEA-NOC/20080201-3 [UPDATE] Multiple flaps on backbone link from Limerick to Cork

HEA-NOC/20080201-3 [UPDATE] Multiple flaps on backbone link from Limerick to Cork

Louis Twomey <[log in to unmask]>

Fri, 1 Feb 2008 16:51:46 +0000

43 lines

HEA-NOC/20080201-3 [UPDATE] Multiple flaps on backbone link from Limerick to Cork

Louis Twomey <[log in to unmask]>

Fri, 1 Feb 2008 12:45:13 +0000

39 lines

New Thread

HEA-NOC/20080201-6 [CLOSE] Scheduled outage on STACS link

HEA-NOC/20080201-6 [CLOSE] Scheduled outage on STACS link

Louis Twomey <[log in to unmask]>

Mon, 25 Feb 2008 09:23:55 +0000

38 lines

New Thread

HEA-NOC/20080201-6 [OPEN] Scheduled outage on STACS link

HEA-NOC/20080201-6 [OPEN] Scheduled outage on STACS link

Louis Twomey <[log in to unmask]>

Fri, 1 Feb 2008 15:56:36 +0000

31 lines

New Thread

HEA-NOC/20080202-1 [OPEN] Commodity IP service maintenance

HEA-NOC/20080202-1 [OPEN] Commodity IP service maintenance

Daniel Lete <[log in to unmask]>

Sat, 2 Feb 2008 08:37:08 +0000

34 lines

New Thread

HEA-NOC/20080202-3 [OPEN] Hermes connectivity loss

HEA-NOC/20080202-3 [OPEN] Hermes connectivity loss

Daniel Lete <[log in to unmask]>

Sat, 2 Feb 2008 22:37:50 +0000

31 lines

New Thread

HEA-NOC/20080202-3 [UPDATE] Hermes connectivity loss

HEA-NOC/20080202-3 [UPDATE] Hermes connectivity loss

Daniel Lete <[log in to unmask]>

Mon, 4 Feb 2008 10:50:41 +0000

46 lines

HEA-NOC/20080202-3 [UPDATE] Hermes connectivity loss

Daniel Lete <[log in to unmask]>

Sun, 3 Feb 2008 13:39:34 +0000

44 lines

HEA-NOC/20080202-3 [UPDATE] Hermes connectivity loss

Daniel Lete <[log in to unmask]>

Sun, 3 Feb 2008 13:02:07 +0000

35 lines

New Thread

HEA-NOC/20080204-2 [CLOSE] St. Angela's Link Instability & Loss of Connectivity

HEA-NOC/20080204-2 [CLOSE] St. Angela's Link Instability & Loss of Connectivity

Brian Nisbet <[log in to unmask]>

Tue, 5 Feb 2008 16:29:18 +0000

50 lines

New Thread

HEA-NOC/20080204-2 [OPEN] St. Angela's LInk Instability & Loss of Connectivity

HEA-NOC/20080204-2 [OPEN] St. Angela's LInk Instability & Loss of Connectivity

Brian Nisbet <[log in to unmask]>

Mon, 4 Feb 2008 10:12:59 +0000

37 lines

New Thread

HEA-NOC/20080204-2 [UPDATE] St. Angela's Link Instability & Loss of Connectivity

HEA-NOC/20080204-2 [UPDATE] St. Angela's Link Instability & Loss of Connectivity

Brian Nisbet <[log in to unmask]>

Mon, 4 Feb 2008 16:19:34 +0000

47 lines

HEA-NOC/20080204-2 [UPDATE] St. Angela's Link Instability & Loss of Connectivity

Brian Nisbet <[log in to unmask]>

Mon, 4 Feb 2008 14:56:40 +0000

42 lines

New Thread

HEA-NOC/20080204-5 [CLOSE] IADT Migration to New IP Equipment

HEA-NOC/20080204-5 [CLOSE] IADT Migration to New IP Equipment

Garwin Liu <[log in to unmask]>

Tue, 12 Feb 2008 14:39:48 +0000

49 lines

New Thread

HEA-NOC/20080204-5 [OPEN] IADT Migration to New IP Equipment

HEA-NOC/20080204-5 [OPEN] IADT Migration to New IP Equipment

Garwin Liu <[log in to unmask]>

Mon, 4 Feb 2008 13:12:50 +0000

41 lines

New Thread

HEA-NOC/20080204-5 [UPDATE] IADT Migration to New IP Equipment

HEA-NOC/20080204-5 [UPDATE] IADT Migration to New IP Equipment

Garwin Liu <[log in to unmask]>

Thu, 7 Feb 2008 09:34:55 +0000

46 lines

New Thread

HEA-NOC/20080204-8 [CLOSE] Loss of Connectivity to IADT & Reslience to NUIM & GEANT

HEA-NOC/20080204-8 [CLOSE] Loss of Connectivity to IADT & Reslience to NUIM & GEANT

Brian Nisbet <[log in to unmask]>

Mon, 11 Feb 2008 09:14:21 +0000

51 lines

New Thread

HEA-NOC/20080204-8 [OPEN] Loss Of Reslience to NUIM

HEA-NOC/20080204-8 [OPEN] Loss Of Reslience to NUIM

Brian Nisbet <[log in to unmask]>

Mon, 4 Feb 2008 16:01:05 +0000

35 lines

New Thread

HEA-NOC/20080204-8 [UPDATE] Loss of Connectivity to IADT & Loss Of Reslience to NUIM

HEA-NOC/20080204-8 [UPDATE] Loss of Connectivity to IADT & Loss Of Reslience to NUIM

Brian Nisbet <[log in to unmask]>

Tue, 5 Feb 2008 10:09:32 +0000

46 lines

New Thread

HEA-NOC/20080204-8 [UPDATE] Loss of Connectivity to IADT & Reslience to NUIM & GEANT

HEA-NOC/20080204-8 [UPDATE] Loss of Connectivity to IADT & Reslience to NUIM & GEANT

Brian Nisbet <[log in to unmask]>

Fri, 8 Feb 2008 20:00:53 +0000

49 lines

New Thread

HEA-NOC/20080205-1 [CLOSE] IPv6 root server addresses

HEA-NOC/20080205-1 [CLOSE] IPv6 root server addresses

Rob Gallagher <[log in to unmask]>

Wed, 6 Feb 2008 09:21:46 +0000

37 lines

New Thread

HEA-NOC/20080205-1 [OPEN] IPv6 root server addresses

HEA-NOC/20080205-1 [OPEN] IPv6 root server addresses

Rob Gallagher <[log in to unmask]>

Tue, 5 Feb 2008 10:24:07 +0000

31 lines

New Thread

HEA-NOC/20080205-1 [UPDATE] IPv6 root server addresses

HEA-NOC/20080205-1 [UPDATE] IPv6 root server addresses

Rob Gallagher <[log in to unmask]>

Tue, 5 Feb 2008 11:06:13 +0000

35 lines

New Thread

HEA-NOC/20080205-13 [OPEN] Geant Primary Link Down

HEA-NOC/20080205-13 [OPEN] Geant Primary Link Down

Brian Nisbet <[log in to unmask]>

Tue, 5 Feb 2008 13:04:59 +0000

32 lines

New Thread

HEA-NOC/20080205-13 [UPDATE] Geant Primary Link Down

HEA-NOC/20080205-13 [UPDATE] Geant Primary Link Down

Brian Nisbet <[log in to unmask]>

Thu, 7 Feb 2008 11:37:38 +0000

48 lines

HEA-NOC/20080205-13 [UPDATE] Geant Primary Link Down

Brian Nisbet <[log in to unmask]>

Thu, 7 Feb 2008 10:07:19 +0000

44 lines

HEA-NOC/20080205-13 [UPDATE] Geant Primary Link Down

Brian Nisbet <[log in to unmask]>

Wed, 6 Feb 2008 08:06:05 +0000

40 lines

HEA-NOC/20080205-13 [UPDATE] Geant Primary Link Down

Brian Nisbet <[log in to unmask]>

Tue, 5 Feb 2008 17:36:20 +0000

36 lines

New Thread

HEA-NOC/20080205-18 [CLOSE] Connectivity to Carlow College lost

HEA-NOC/20080205-18 [CLOSE] Connectivity to Carlow College lost

Rob Gallagher <[log in to unmask]>

Wed, 6 Feb 2008 13:58:32 +0000

41 lines

New Thread

HEA-NOC/20080205-18 [OPEN] Connectivity to Carlow College lost

HEA-NOC/20080205-18 [OPEN] Connectivity to Carlow College lost

Rob Gallagher <[log in to unmask]>

Tue, 5 Feb 2008 15:07:32 +0000

30 lines

New Thread

HEA-NOC/20080205-18 [UPDATE] Connectivity to Carlow College lost

HEA-NOC/20080205-18 [UPDATE] Connectivity to Carlow College lost

Rob Gallagher <[log in to unmask]>

Tue, 5 Feb 2008 15:47:49 +0000

35 lines

New Thread

HEA-NOC/20080205-2 [CLOSE] Temporary Loss of Routing to Citywest Core Router

HEA-NOC/20080205-2 [CLOSE] Temporary Loss of Routing to Citywest Core Router

Brian Nisbet <[log in to unmask]>

Wed, 6 Feb 2008 10:15:44 +0000

45 lines

New Thread

HEA-NOC/20080205-2 [OPEN] Temporary Loss of Routing to Citywest Core Router

HEA-NOC/20080205-2 [OPEN] Temporary Loss of Routing to Citywest Core Router

Brian Nisbet <[log in to unmask]>

Tue, 5 Feb 2008 10:46:09 +0000

37 lines

New Thread

HEA-NOC/20080205-2 [UPDATE] Temporary Loss of Routing to Citywest Core Router

HEA-NOC/20080205-2 [UPDATE] Temporary Loss of Routing to Citywest Core Router

Gareth Eason <[log in to unmask]>

Tue, 5 Feb 2008 13:14:05 +0000

43 lines

HEA-NOC/20080205-2 [UPDATE] Temporary Loss of Routing to Citywest Core Router

Gareth Eason <[log in to unmask]>

Tue, 5 Feb 2008 12:58:47 +0000

40 lines

New Thread

HEA-NOC/20080205-8 [CLOSE] Package updates and kernel upgrades on multiple machines

HEA-NOC/20080205-8 [CLOSE] Package updates and kernel upgrades on multiple machines

Gareth Eason <[log in to unmask]>

Tue, 19 Feb 2008 13:19:09 +0000

66 lines

New Thread

HEA-NOC/20080205-8 [OPEN] Package updates and kernel upgrades on multiple machines

HEA-NOC/20080205-8 [OPEN] Package updates and kernel upgrades on multiple machines

Rob Gallagher <[log in to unmask]>

Tue, 5 Feb 2008 12:05:50 +0000

42 lines

New Thread

HEA-NOC/20080205-8 [UPDATE] Package updates and kernel upgrades on multiple machines

HEA-NOC/20080205-8 [UPDATE] Package updates and kernel upgrades on multiple machines

Gareth Eason <[log in to unmask]>

Mon, 18 Feb 2008 18:12:36 +0000

61 lines

HEA-NOC/20080205-8 [UPDATE] Package updates and kernel upgrades on multiple machines

Rob Gallagher <[log in to unmask]>

Fri, 15 Feb 2008 14:47:09 +0000

56 lines

HEA-NOC/20080205-8 [UPDATE] Package updates and kernel upgrades on multiple machines

Rob Gallagher <[log in to unmask]>

Wed, 13 Feb 2008 10:18:52 +0000

54 lines

HEA-NOC/20080205-8 [UPDATE] Package updates and kernel upgrades on multiple machines

Rob Gallagher <[log in to unmask]>

Mon, 11 Feb 2008 12:12:37 +0000

49 lines

HEA-NOC/20080205-8 [UPDATE] Package updates and kernel upgrades on multiple machines

Louis Twomey <[log in to unmask]>

Wed, 6 Feb 2008 10:08:53 +0000

47 lines

New Thread

HEA-NOC/20080206-10 [OPEN] Multiple flaps on HRB link

HEA-NOC/20080206-10 [OPEN] Multiple flaps on HRB link

Rob Gallagher <[log in to unmask]>

Wed, 6 Feb 2008 14:07:34 +0000

28 lines

New Thread

HEA-NOC/20080206-10 [UPDATE] Multiple flaps on HRB link

HEA-NOC/20080206-10 [UPDATE] Multiple flaps on HRB link

Rob Gallagher <[log in to unmask]>

Wed, 6 Feb 2008 17:20:52 +0000

35 lines

New Thread

HEA-NOC/20080206-14 [OPEN/CLOSE] HEAnet mail server problem

HEA-NOC/20080206-14 [OPEN/CLOSE] HEAnet mail server problem

Cillian Sharkey <[log in to unmask]>

Wed, 6 Feb 2008 16:59:21 +0000

32 lines

New Thread

HEA-NOC/20080206-3 [CLOSE] Loss of Connection to IADT

HEA-NOC/20080206-3 [CLOSE] Loss of Connection to IADT

Brian Nisbet <[log in to unmask]>

Wed, 6 Feb 2008 17:30:53 +0000

39 lines

New Thread

HEA-NOC/20080206-3 [OPEN] Loss of Connection to IADT

HEA-NOC/20080206-3 [OPEN] Loss of Connection to IADT

Brian Nisbet <[log in to unmask]>

Wed, 6 Feb 2008 09:53:32 +0000

32 lines

New Thread

HEA-NOC/20080206-3 [UPDATE] Loss of Connection to IADT

HEA-NOC/20080206-3 [UPDATE] Loss of Connection to IADT

Brian Nisbet <[log in to unmask]>

Wed, 6 Feb 2008 10:15:04 +0000

37 lines

New Thread

HEA-NOC/20080206-4 [OPEN] Enabling Netflow on CRS

HEA-NOC/20080206-4 [OPEN] Enabling Netflow on CRS

Cormac O'Cianain <[log in to unmask]>

Wed, 6 Feb 2008 10:23:06 +0000

31 lines

New Thread

HEA-NOC/20080206-4 [UPDATE] Enabling Netflow on CRS

HEA-NOC/20080206-4 [UPDATE] Enabling Netflow on CRS

Cormac O'Cianain <[log in to unmask]>

Wed, 13 Feb 2008 08:42:43 +0000

45 lines

HEA-NOC/20080206-4 [UPDATE] Enabling Netflow on CRS

Cormac O'Cianain <[log in to unmask]>

Thu, 7 Feb 2008 10:45:18 +0000

42 lines

HEA-NOC/20080206-4 [UPDATE] Enabling Netflow on CRS

Cormac O'Cianain <[log in to unmask]>

Wed, 6 Feb 2008 16:46:45 +0000

38 lines

HEA-NOC/20080206-4 [UPDATE] Enabling Netflow on CRS

Cormac O'Cianain <[log in to unmask]>

Wed, 6 Feb 2008 13:31:01 +0000

35 lines

New Thread

HEA-NOC/20080207-1 [CLOSE] Loss of Connection to St Angela's

HEA-NOC/20080207-1 [CLOSE] Loss of Connection to St Angela's

Brian Nisbet <[log in to unmask]>

Fri, 8 Feb 2008 20:01:17 +0000

43 lines

New Thread

HEA-NOC/20080207-1 [OPEN] Loss of Connection to St Angela's

HEA-NOC/20080207-1 [OPEN] Loss of Connection to St Angela's

Brian Nisbet <[log in to unmask]>

Thu, 7 Feb 2008 09:23:41 +0000

32 lines

New Thread

HEA-NOC/20080207-1 [UPDATE] Loss of Connection to St Angela's

HEA-NOC/20080207-1 [UPDATE] Loss of Connection to St Angela's

Brian Nisbet <[log in to unmask]>

Thu, 7 Feb 2008 16:08:35 +0000

41 lines

HEA-NOC/20080207-1 [UPDATE] Loss of Connection to St Angela's

Brian Nisbet <[log in to unmask]>

Thu, 7 Feb 2008 11:18:45 +0000

37 lines

New Thread

HEA-NOC/20080207-11 [CLOSE] Loss of signal on adva equipment between Kilcarbery and Citywest

HEA-NOC/20080207-11 [CLOSE] Loss of signal on adva equipment between Kilcarbery and Citywest

Rob Gallagher <[log in to unmask]>

Thu, 7 Feb 2008 15:38:27 +0000

33 lines

New Thread

HEA-NOC/20080207-11 [OPEN] Loss of signal on adva equipment between Kilcarbery and Citywest

HEA-NOC/20080207-11 [OPEN] Loss of signal on adva equipment between Kilcarbery and Citywest

Rob Gallagher <[log in to unmask]>

Thu, 7 Feb 2008 15:31:46 +0000

29 lines

New Thread

HEA-NOC/20080207-12 [OPEN] GÉANT replacement STM-16

HEA-NOC/20080207-12 [OPEN] GÉANT replacement STM-16

Dave Wilson <[log in to unmask]>

Thu, 7 Feb 2008 17:13:52 +0000

43 lines

New Thread

HEA-NOC/20080207-12 [UPDATE] GÉANT replacement STM-16

HEA-NOC/20080207-12 [UPDATE] GÉANT replacement STM-16

Dave Wilson <[log in to unmask]>

Thu, 21 Feb 2008 12:52:33 +0000

64 lines

HEA-NOC/20080207-12 [UPDATE] GÉANT replacement STM-16

Dave Wilson <[log in to unmask]>

Wed, 20 Feb 2008 13:38:19 +0000

64 lines

HEA-NOC/20080207-12 [UPDATE] GÉANT replacement STM-16

Dave Wilson <[log in to unmask]>

Wed, 13 Feb 2008 17:14:39 +0000

59 lines

HEA-NOC/20080207-12 [UPDATE] GÉANT replacement STM-16

Dave Wilson <[log in to unmask]>

Wed, 13 Feb 2008 12:38:28 +0000

55 lines

HEA-NOC/20080207-12 [UPDATE] GÉANT replacement STM-16

Dave Wilson <[log in to unmask]>

Tue, 12 Feb 2008 06:33:52 +0000

52 lines

HEA-NOC/20080207-12 [UPDATE] GÉANT replacement STM-16

Dave Wilson <[log in to unmask]>

Fri, 8 Feb 2008 17:12:18 +0000

49 lines

New Thread

HEA-NOC/20080207-2 [CLOSE] Power Failure in TCD PoP

HEA-NOC/20080207-2 [CLOSE] Power Failure in TCD PoP

Brian Nisbet <[log in to unmask]>

Fri, 8 Feb 2008 12:59:59 +0000

41 lines

New Thread

HEA-NOC/20080207-2 [OPEN] Power Failure in TCD PoP

HEA-NOC/20080207-2 [OPEN] Power Failure in TCD PoP

Brian Nisbet <[log in to unmask]>

Thu, 7 Feb 2008 09:54:22 +0000

36 lines

New Thread

HEA-NOC/20080208-11 [CLOSE] Loss of Geant Secondary BGP Peering

HEA-NOC/20080208-11 [CLOSE] Loss of Geant Secondary BGP Peering

Brian Nisbet <[log in to unmask]>

Mon, 11 Feb 2008 09:12:15 +0000

38 lines

New Thread

HEA-NOC/20080208-11 [OPEN] Loss of Geant Secondary BGP Peering

HEA-NOC/20080208-11 [OPEN] Loss of Geant Secondary BGP Peering

Brian Nisbet <[log in to unmask]>

Fri, 8 Feb 2008 16:45:57 +0000

33 lines

New Thread

HEA-NOC/20080208-13 [CLOSE] Unable to divert NOC number to answering service

HEA-NOC/20080208-13 [CLOSE] Unable to divert NOC number to answering service

Gareth Eason <[log in to unmask]>

Tue, 12 Feb 2008 13:12:27 +0000

40 lines

New Thread

HEA-NOC/20080208-13 [OPEN] Unable to divert NOC number to answering service

HEA-NOC/20080208-13 [OPEN] Unable to divert NOC number to answering service

Gareth Eason <[log in to unmask]>

Fri, 8 Feb 2008 20:55:24 +0000

34 lines

New Thread

HEA-NOC/20080208-2 [CLOSE] Transition CIT To New IP Network

HEA-NOC/20080208-2 [CLOSE] Transition CIT To New IP Network

Brian Nisbet <[log in to unmask]>

Wed, 13 Feb 2008 10:15:48 +0000

49 lines

New Thread

HEA-NOC/20080208-2 [OPEN] Transition CIT To New IP Network

HEA-NOC/20080208-2 [OPEN] Transition CIT To New IP Network

Brian Nisbet <[log in to unmask]>

Fri, 8 Feb 2008 10:08:52 +0000

38 lines

New Thread

HEA-NOC/20080208-2 [UPDATE] Transition CIT To New IP Network

HEA-NOC/20080208-2 [UPDATE] Transition CIT To New IP Network

Brian Nisbet <[log in to unmask]>

Tue, 12 Feb 2008 08:44:29 +0000

46 lines

HEA-NOC/20080208-2 [UPDATE] Transition CIT To New IP Network

Brian Nisbet <[log in to unmask]>

Mon, 11 Feb 2008 16:03:03 +0000

42 lines

New Thread

HEA-NOC/20080208-4 [CLOSE] Cessation of legacy ITnet Commodity IP service

HEA-NOC/20080208-4 [CLOSE] Cessation of legacy ITnet Commodity IP service

Owen Byrne <[log in to unmask]>

Mon, 25 Feb 2008 15:19:38 +0000

45 lines

New Thread

HEA-NOC/20080208-4 [OPEN] Cessation of legacy ITnet Commodity IP service

HEA-NOC/20080208-4 [OPEN] Cessation of legacy ITnet Commodity IP service

Owen Byrne <[log in to unmask]>

Fri, 8 Feb 2008 11:15:51 +0000

32 lines

New Thread

HEA-NOC/20080208-4 [UPDATE] Cessation of legacy ITnet Commodity IP service

HEA-NOC/20080208-4 [UPDATE] Cessation of legacy ITnet Commodity IP service

Owen Byrne <[log in to unmask]>

Fri, 8 Feb 2008 12:18:59 +0000

38 lines

New Thread

HEA-NOC/20080209-1 [CLOSE] Connectivity issues to UL

HEA-NOC/20080209-1 [CLOSE] Connectivity issues to UL

Rachael Holt <[log in to unmask]>

Mon, 11 Feb 2008 09:22:53 +0000

59 lines

New Thread

HEA-NOC/20080209-1 [UPDATE] Connectivity issues to UL

HEA-NOC/20080209-1 [UPDATE] Connectivity issues to UL

Rachael Holt <[log in to unmask]>

Sun, 10 Feb 2008 15:12:01 +0000

52 lines

New Thread

HEA-NOC/20080209-3 [CLOSE] NUI, Primary Link Down

HEA-NOC/20080209-3 [CLOSE] NUI, Primary Link Down

Rachael Holt <[log in to unmask]>

Mon, 11 Feb 2008 14:27:33 +0000

40 lines

New Thread

HEA-NOC/20080209-3 [OPEN] NUI, Primary Link Down

HEA-NOC/20080209-3 [OPEN] NUI, Primary Link Down

Rachael Holt <[log in to unmask]>

Sat, 9 Feb 2008 16:52:27 +0000

27 lines

New Thread

HEA-NOC/20080209-3 [UPDATE] NUI, Primary Link Down

HEA-NOC/20080209-3 [UPDATE] NUI, Primary Link Down

Rachael Holt <[log in to unmask]>

Mon, 11 Feb 2008 09:27:50 +0000

36 lines

HEA-NOC/20080209-3 [UPDATE] NUI, Primary Link Down

Rachael Holt <[log in to unmask]>

Sun, 10 Feb 2008 16:58:15 +0000

32 lines

HEA-NOC/20080209-3 [UPDATE] NUI, Primary Link Down

Rachael Holt <[log in to unmask]>

Sun, 10 Feb 2008 10:51:36 +0000

30 lines

New Thread

HEA-NOC/20080209-5 [OPEN/CLOSE] Loss of connectivity to DIAS

HEA-NOC/20080209-5 [OPEN/CLOSE] Loss of connectivity to DIAS

Rachael Holt <[log in to unmask]>

Sat, 9 Feb 2008 21:02:33 +0000

48 lines

New Thread

HEA-NOC/20080209-6 [CLOSE] Connectivity issues to IUA

HEA-NOC/20080209-6 [CLOSE] Connectivity issues to IUA

Brian Boyle <[log in to unmask]>

Wed, 13 Feb 2008 16:37:49 +0000

49 lines

New Thread

HEA-NOC/20080209-6 [OPEN] Connectivity issues to IUA

HEA-NOC/20080209-6 [OPEN] Connectivity issues to IUA

Rachael Holt <[log in to unmask]>

Sat, 9 Feb 2008 21:11:11 +0000

31 lines

New Thread

HEA-NOC/20080209-6 [UPDATE] Connectivity issues to IUA

HEA-NOC/20080209-6 [UPDATE] Connectivity issues to IUA

Brian Boyle <[log in to unmask]>

Mon, 11 Feb 2008 09:47:44 +0000

44 lines

HEA-NOC/20080209-6 [UPDATE] Connectivity issues to IUA

Brian Boyle <[log in to unmask]>

Mon, 11 Feb 2008 09:37:56 +0000

40 lines

HEA-NOC/20080209-6 [UPDATE] Connectivity issues to IUA

Rachael Holt <[log in to unmask]>

Mon, 11 Feb 2008 09:35:35 +0000

37 lines

HEA-NOC/20080209-6 [UPDATE] Connectivity issues to IUA

Brian Boyle <[log in to unmask]>

Mon, 11 Feb 2008 09:34:10 +0000

35 lines

New Thread

HEA-NOC/20080209-7 [CLOSE] Connectivity issues to BCFE

HEA-NOC/20080209-7 [CLOSE] Connectivity issues to BCFE

Rachael Holt <[log in to unmask]>

Mon, 11 Feb 2008 10:15:07 +0000

40 lines

New Thread

HEA-NOC/20080209-7 [OPEN] Connectivity issues to BCFE

HEA-NOC/20080209-7 [OPEN] Connectivity issues to BCFE

Rachael Holt <[log in to unmask]>

Sat, 9 Feb 2008 21:17:31 +0000

31 lines

New Thread

HEA-NOC/20080210-2 [OPEN] No connectivity to cdveckillester1

HEA-NOC/20080210-2 [OPEN] No connectivity to cdveckillester1

Rachael Holt <[log in to unmask]>

Sun, 10 Feb 2008 11:42:55 +0000

29 lines

New Thread

HEA-NOC/20080210-2 [UPDATE] No connectivity to cdveckillester1

HEA-NOC/20080210-2 [UPDATE] No connectivity to cdveckillester1

Daniel Lete <[log in to unmask]>

Tue, 26 Feb 2008 16:37:05 +0000

48 lines

HEA-NOC/20080210-2 [UPDATE] No connectivity to cdveckillester1

Gareth Eason <[log in to unmask]>

Mon, 18 Feb 2008 12:45:50 +0000

46 lines

HEA-NOC/20080210-2 [UPDATE] No connectivity to cdveckillester1

Rob Gallagher <[log in to unmask]>

Mon, 18 Feb 2008 09:43:34 +0000

42 lines

HEA-NOC/20080210-2 [UPDATE] No connectivity to cdveckillester1

Paul Cosgrove <[log in to unmask]>

Tue, 12 Feb 2008 13:02:12 +0000

36 lines

HEA-NOC/20080210-2 [UPDATE] No connectivity to cdveckillester1

Brian Boyle <[log in to unmask]>

Tue, 12 Feb 2008 10:07:33 +0000

33 lines

HEA-NOC/20080210-2 [UPDATE] No connectivity to cdveckillester1

Rachael Holt <[log in to unmask]>

Mon, 11 Feb 2008 09:30:55 +0000

30 lines

New Thread

HEA-NOC/20080211-6 [CLOSE] Installation of 10G cards on all INEX switchs

HEA-NOC/20080211-6 [CLOSE] Installation of 10G cards on all INEX switchs

Brian Boyle <[log in to unmask]>

Tue, 26 Feb 2008 15:46:15 +0000

45 lines

New Thread

HEA-NOC/20080211-6 [OPEN] Installation of 10G cards on all INEX switchs

HEA-NOC/20080211-6 [OPEN] Installation of 10G cards on all INEX switchs

Brian Boyle <[log in to unmask]>

Mon, 11 Feb 2008 15:02:59 +0000

35 lines

New Thread

HEA-NOC/20080212-11 [OPEN] Report of slow connectivity from Cork School of Music

HEA-NOC/20080212-11 [OPEN] Report of slow connectivity from Cork School of Music

Paul Cosgrove <[log in to unmask]>

Tue, 12 Feb 2008 16:46:03 +0000

32 lines

New Thread

HEA-NOC/20080212-11 [UPDATE] Report of slow connectivity from Cork School of Music

HEA-NOC/20080212-11 [UPDATE] Report of slow connectivity from Cork School of Music

Paul Cosgrove <[log in to unmask]>

Fri, 15 Feb 2008 18:10:04 +0000

49 lines

New Thread

HEA-NOC/20080212-2 [CLOSE] Test listserv failover

HEA-NOC/20080212-2 [CLOSE] Test listserv failover

Rob Gallagher <[log in to unmask]>

Thu, 14 Feb 2008 11:30:55 +0000

43 lines

New Thread

HEA-NOC/20080212-2 [OPEN] Test listserv failover

HEA-NOC/20080212-2 [OPEN] Test listserv failover

Rob Gallagher <[log in to unmask]>

Tue, 12 Feb 2008 11:52:49 +0000

34 lines

New Thread

HEA-NOC/20080212-2 [UPDATE] Test listserv failover

HEA-NOC/20080212-2 [UPDATE] Test listserv failover

Rob Gallagher <[log in to unmask]>

Tue, 12 Feb 2008 17:49:36 +0000

41 lines

HEA-NOC/20080212-2 [UPDATE] Test listserv failover

Rob Gallagher <[log in to unmask]>

Tue, 12 Feb 2008 17:32:26 +0000

37 lines

New Thread

HEA-NOC/20080212-4 [CLOSE] Transition LIT To New IP Network

HEA-NOC/20080212-4 [CLOSE] Transition LIT To New IP Network

Tony Gray <[log in to unmask]>

Wed, 27 Feb 2008 15:48:58 +0000

45 lines

New Thread

HEA-NOC/20080212-4 [OPEN] Transition LIT To New IP Network

HEA-NOC/20080212-4 [OPEN] Transition LIT To New IP Network

Tony Gray <[log in to unmask]>

Tue, 12 Feb 2008 12:21:07 +0000

37 lines

New Thread

HEA-NOC/20080212-4 [UPDATE] Transition LIT To New IP Network

HEA-NOC/20080212-4 [UPDATE] Transition LIT To New IP Network

Tony Gray <[log in to unmask]>

Thu, 14 Feb 2008 09:03:34 +0000

43 lines

HEA-NOC/20080212-4 [UPDATE] Transition LIT To New IP Network

Tony Gray <[log in to unmask]>

Tue, 12 Feb 2008 17:38:47 +0000

40 lines

New Thread

HEA-NOC/20080212-6 [CLOSE] Loss of connectivity to GDA

HEA-NOC/20080212-6 [CLOSE] Loss of connectivity to GDA

Paul Cosgrove <[log in to unmask]>

Tue, 12 Feb 2008 13:24:14 +0000

35 lines

New Thread

HEA-NOC/20080212-6 [OPEN] Loss of connectivity to GDA

HEA-NOC/20080212-6 [OPEN] Loss of connectivity to GDA

Paul Cosgrove <[log in to unmask]>

Tue, 12 Feb 2008 13:22:27 +0000

27 lines

New Thread

HEA-NOC/20080213-1 [CLOSE] The servers hosting the video streaming service are down

HEA-NOC/20080213-1 [CLOSE] The servers hosting the video streaming service are down

Brian Boyle <[log in to unmask]>

Tue, 19 Feb 2008 15:15:52 +0000

38 lines

New Thread

HEA-NOC/20080213-1 [OPEN] The servers hosting the video streaming service are down

HEA-NOC/20080213-1 [OPEN] The servers hosting the video streaming service are down

Brian Boyle <[log in to unmask]>

Wed, 13 Feb 2008 09:19:35 +0000

29 lines

New Thread

HEA-NOC/20080213-1 [UPDATE] The servers hosting the video streaming service are down

HEA-NOC/20080213-1 [UPDATE] The servers hosting the video streaming service are down

Brian Boyle <[log in to unmask]>

Wed, 13 Feb 2008 14:29:44 +0000

32 lines

New Thread

HEA-NOC/20080213-3 [OPEN] ESBT Plan to carry out repairs in the Kilbarry/Killonan POPs

HEA-NOC/20080213-3 [OPEN] ESBT Plan to carry out repairs in the Kilbarry/Killonan POPs

Brian Boyle <[log in to unmask]>

Wed, 13 Feb 2008 10:52:50 +0000

32 lines

New Thread

HEA-NOC/20080213-4 [CLOSE] Global Crossing Maintenance

HEA-NOC/20080213-4 [CLOSE] Global Crossing Maintenance

Daniel Lete <[log in to unmask]>

Wed, 27 Feb 2008 08:49:39 +0000

36 lines

New Thread

HEA-NOC/20080213-4 [OPEN] Outage to replace router hardware

HEA-NOC/20080213-4 [OPEN] Outage to replace router hardware

Brian Boyle <[log in to unmask]>

Wed, 13 Feb 2008 11:02:58 +0000

30 lines

New Thread

HEA-NOC/20080213-6 [CLOSE] IT Carlow Commodity IP Service Maintenance

HEA-NOC/20080213-6 [CLOSE] IT Carlow Commodity IP Service Maintenance

Brian Nisbet <[log in to unmask]>

Wed, 27 Feb 2008 08:42:02 +0000

38 lines

New Thread

HEA-NOC/20080213-6 [OPEN] IT Carlow Commodity IP Service Maintenance

HEA-NOC/20080213-6 [OPEN] IT Carlow Commodity IP Service Maintenance

Brian Nisbet <[log in to unmask]>

Wed, 13 Feb 2008 15:10:45 +0000

35 lines

New Thread

HEA-NOC/20080214-9 [OPEN] ESB Planned Maintenance affecting Kilbarry/Killoteran Fibre

HEA-NOC/20080214-9 [OPEN] ESB Planned Maintenance affecting Kilbarry/Killoteran Fibre

Paul Cosgrove <[log in to unmask]>

Thu, 14 Feb 2008 13:12:11 +0000

46 lines

New Thread

HEA-NOC/20080215-1 [OPEN] Fault repair on fibres in the Claremorris-Ballina duct

HEA-NOC/20080215-1 [OPEN] Fault repair on fibres in the Claremorris-Ballina duct

Brian Boyle <[log in to unmask]>

Fri, 15 Feb 2008 09:59:21 +0000

31 lines

New Thread

HEA-NOC/20080215-3 [OPEN] ESBT Plan to carry out repairs in the Kilbarry/Killoteran POPs

HEA-NOC/20080215-3 [OPEN] ESBT Plan to carry out repairs in the Kilbarry/Killoteran POPs

Brian Boyle <[log in to unmask]>

Fri, 15 Feb 2008 10:03:31 +0000

31 lines

New Thread

HEA-NOC/20080215-6 [CLOSE] Listserv and SixXS6 server upgrade

HEA-NOC/20080215-6 [CLOSE] Listserv and SixXS6 server upgrade

Aidan Carty <[log in to unmask]>

Wed, 27 Feb 2008 10:55:26 +0000

40 lines

New Thread

HEA-NOC/20080215-6 [OPEN] Listserv and SixXS6 server upgrade

HEA-NOC/20080215-6 [OPEN] Listserv and SixXS6 server upgrade

Aidan Carty <[log in to unmask]>

Fri, 15 Feb 2008 11:22:22 +0000

32 lines

New Thread

HEA-NOC/20080215-6 [UPDATE] Listserv and SixXS6 server upgrade

HEA-NOC/20080215-6 [UPDATE] Listserv and SixXS6 server upgrade

Aidan Carty <[log in to unmask]>

Thu, 21 Feb 2008 15:37:12 +0000

38 lines

HEA-NOC/20080215-6 [UPDATE] Listserv and SixXS6 server upgrade

Aidan Carty <[log in to unmask]>

Wed, 20 Feb 2008 14:59:25 +0000

35 lines

New Thread

HEA-NOC/20080215-8 [OPEN] Scheduled Power Outage at CIT on 22nd Feb

HEA-NOC/20080215-8 [OPEN] Scheduled Power Outage at CIT on 22nd Feb

Paul Cosgrove <[log in to unmask]>

Fri, 15 Feb 2008 14:42:55 +0000

40 lines

New Thread

HEA-NOC/20080218-8 [CLOSE] Planned power outage in WIT Carriganore

HEA-NOC/20080218-8 [CLOSE] Planned power outage in WIT Carriganore

Daniel Lete <[log in to unmask]>

Wed, 27 Feb 2008 08:47:10 +0000

45 lines

New Thread

HEA-NOC/20080218-8 [OPEN] Planned power outage in WIT Carriganore

HEA-NOC/20080218-8 [OPEN] Planned power outage in WIT Carriganore

Gareth Eason <[log in to unmask]>

Mon, 18 Feb 2008 16:13:24 +0000

35 lines

New Thread

HEA-NOC/20080218-8 [UPDATE] Planned power outage in WIT Carriganore

HEA-NOC/20080218-8 [UPDATE] Planned power outage in WIT Carriganore

Louis Twomey <[log in to unmask]>

Tue, 26 Feb 2008 09:39:22 +0000

43 lines

HEA-NOC/20080218-8 [UPDATE] Planned power outage in WIT Carriganore

Daniel Lete <[log in to unmask]>

Mon, 25 Feb 2008 14:46:30 +0000

40 lines

New Thread

HEA-NOC/20080219-2 [CLOSE] Circuits down to VEC Killester and Liberties

HEA-NOC/20080219-2 [CLOSE] Circuits down to VEC Killester and Liberties

Gareth Eason <[log in to unmask]>

Tue, 19 Feb 2008 15:53:29 +0000

56 lines

New Thread

HEA-NOC/20080219-2 [OPEN] Circuits down to VEC Killester and Liberties

HEA-NOC/20080219-2 [OPEN] Circuits down to VEC Killester and Liberties

Gareth Eason <[log in to unmask]>

Tue, 19 Feb 2008 10:36:37 +0000

34 lines

New Thread

HEA-NOC/20080219-2 [UPDATE] Circuits down to VEC Killester and Liberties

HEA-NOC/20080219-2 [UPDATE] Circuits down to VEC Killester and Liberties

Gareth Eason <[log in to unmask]>

Tue, 19 Feb 2008 12:46:52 +0000

49 lines

HEA-NOC/20080219-2 [UPDATE] Circuits down to VEC Killester and Liberties

Gareth Eason <[log in to unmask]>

Tue, 19 Feb 2008 10:41:03 +0000

37 lines

New Thread

HEA-NOC/20080219-3 [CLOSE] Hardware failure of samhain, web services server

HEA-NOC/20080219-3 [CLOSE] Hardware failure of samhain, web services server

Gareth Eason <[log in to unmask]>

Wed, 20 Feb 2008 12:32:36 +0000

74 lines

New Thread

HEA-NOC/20080219-3 [OPEN] Hardware failure of samhain, web services server

HEA-NOC/20080219-3 [OPEN] Hardware failure of samhain, web services server

Gareth Eason <[log in to unmask]>

Tue, 19 Feb 2008 10:58:20 +0000

52 lines

New Thread

HEA-NOC/20080219-3 [UPDATE] Hardware failure of samhain, web services server

HEA-NOC/20080219-3 [UPDATE] Hardware failure of samhain, web services server

Gareth Eason <[log in to unmask]>

Tue, 19 Feb 2008 13:12:01 +0000

71 lines

HEA-NOC/20080219-3 [UPDATE] Hardware failure of samhain, web services server

Gareth Eason <[log in to unmask]>

Tue, 19 Feb 2008 12:07:35 +0000

62 lines

New Thread

HEA-NOC/20080221-1 [OPEN] ATM Interface to St. Angela's flapping

HEA-NOC/20080221-1 [OPEN] ATM Interface to St. Angela's flapping

Gareth Eason <[log in to unmask]>

Thu, 21 Feb 2008 09:44:12 +0000

31 lines

New Thread

HEA-NOC/20080221-1 [UPDATE] ATM Interface to St. Angela's flapping

HEA-NOC/20080221-1 [UPDATE] ATM Interface to St. Angela's flapping

Daniel Lete <[log in to unmask]>

Mon, 25 Feb 2008 12:36:18 +0000

84 lines

HEA-NOC/20080221-1 [UPDATE] ATM Interface to St. Angela's flapping

Daniel Lete <[log in to unmask]>

Mon, 25 Feb 2008 10:26:11 +0000

81 lines

HEA-NOC/20080221-1 [UPDATE] ATM Interface to St. Angela's flapping

Daniel Lete <[log in to unmask]>

Mon, 25 Feb 2008 09:39:28 +0000

72 lines

HEA-NOC/20080221-1 [UPDATE] ATM Interface to St. Angela's flapping

Brian Boyle <[log in to unmask]>

Sun, 24 Feb 2008 12:18:58 +0000

68 lines

HEA-NOC/20080221-1 [UPDATE] ATM Interface to St. Angela's flapping

Brian Nisbet <[log in to unmask]>

Fri, 22 Feb 2008 18:17:08 +0000

65 lines

HEA-NOC/20080221-1 [UPDATE] ATM Interface to St. Angela's flapping

Dave Wilson <[log in to unmask]>

Fri, 22 Feb 2008 16:48:42 +0000

58 lines

HEA-NOC/20080221-1 [UPDATE] ATM Interface to St. Angela's flapping

Dave Wilson <[log in to unmask]>

Fri, 22 Feb 2008 12:32:46 +0000

54 lines

HEA-NOC/20080221-1 [UPDATE] ATM Interface to St. Angela's flapping

Dave Wilson <[log in to unmask]>

Fri, 22 Feb 2008 10:00:05 +0000

51 lines

HEA-NOC/20080221-1 [UPDATE] ATM Interface to St. Angela's flapping

Gareth Eason <[log in to unmask]>

Thu, 21 Feb 2008 10:04:55 +0000

43 lines

New Thread

HEA-NOC/20080221-1 [UPDATE] STACS flapping circuit

HEA-NOC/20080221-1 [UPDATE] STACS flapping circuit

Daniel Lete <[log in to unmask]>

Tue, 26 Feb 2008 12:30:17 +0000

109 lines

HEA-NOC/20080221-1 [UPDATE] STACS flapping circuit

Daniel Lete <[log in to unmask]>

Tue, 26 Feb 2008 11:51:34 +0000

105 lines

HEA-NOC/20080221-1 [UPDATE] STACS flapping circuit

Daniel Lete <[log in to unmask]>

Tue, 26 Feb 2008 11:25:25 +0000

100 lines

HEA-NOC/20080221-1 [UPDATE] STACS flapping circuit

Daniel Lete <[log in to unmask]>

Tue, 26 Feb 2008 09:24:22 +0000

95 lines

HEA-NOC/20080221-1 [UPDATE] STACS flapping circuit

Daniel Lete <[log in to unmask]>

Mon, 25 Feb 2008 17:35:10 +0000

89 lines

New Thread

HEA-NOC/20080221-16 [OPEN] BT fibre repair works at Manuella Junction

HEA-NOC/20080221-16 [OPEN] BT fibre repair works at Manuella Junction

Gareth Eason <[log in to unmask]>

Fri, 22 Feb 2008 09:39:12 +0000

31 lines

New Thread

HEA-NOC/20080221-8 [CLOSE] Configuration change on NCAD CPE to accommodate new network.

HEA-NOC/20080221-8 [CLOSE] Configuration change on NCAD CPE to accommodate new network.

Gareth Eason <[log in to unmask]>

Fri, 22 Feb 2008 13:17:30 +0000

47 lines

New Thread

HEA-NOC/20080221-8 [OPEN] Configuration change on NCAD CPE to accommodate new network.

HEA-NOC/20080221-8 [OPEN] Configuration change on NCAD CPE to accommodate new network.

Gareth Eason <[log in to unmask]>

Thu, 21 Feb 2008 11:14:56 +0000

30 lines

New Thread

HEA-NOC/20080221-8 [UPDATE] Configuration change on NCAD CPE to accommodate new network.

HEA-NOC/20080221-8 [UPDATE] Configuration change on NCAD CPE to accommodate new network.

Gareth Eason <[log in to unmask]>

Thu, 21 Feb 2008 18:23:36 +0000

40 lines

HEA-NOC/20080221-8 [UPDATE] Configuration change on NCAD CPE to accommodate new network.

Gareth Eason <[log in to unmask]>

Thu, 21 Feb 2008 17:58:20 +0000

37 lines

New Thread

HEA-NOC/20080223-1 [CLOSE] The network access swich in the UL Outreach centre rebooted

HEA-NOC/20080223-1 [CLOSE] The network access swich in the UL Outreach centre rebooted

Brian Boyle <[log in to unmask]>

Mon, 25 Feb 2008 17:20:39 +0000

36 lines

New Thread

HEA-NOC/20080223-1 [OPEN] The network access swich in the UL Outreach centre rebooted

HEA-NOC/20080223-1 [OPEN] The network access swich in the UL Outreach centre rebooted

Brian Boyle <[log in to unmask]>

Sat, 23 Feb 2008 12:19:34 +0000

29 lines

New Thread

HEA-NOC/20080223-2 [CLOSE] A circuit to Mary Immaculate college transitioned briefly

HEA-NOC/20080223-2 [CLOSE] A circuit to Mary Immaculate college transitioned briefly

Brian Boyle <[log in to unmask]>

Mon, 25 Feb 2008 17:21:54 +0000

36 lines

New Thread

HEA-NOC/20080223-2 [OPEN] A circuit to Mary Immaculate college transitioned briefly

HEA-NOC/20080223-2 [OPEN] A circuit to Mary Immaculate college transitioned briefly

Brian Boyle <[log in to unmask]>

Sat, 23 Feb 2008 13:50:21 +0000

29 lines

New Thread

HEA-NOC/20080224-1 [CLOSE] Leased lines to two clients in the Mount Street area are down

HEA-NOC/20080224-1 [CLOSE] Leased lines to two clients in the Mount Street area are down

Brian Boyle <[log in to unmask]>

Mon, 25 Feb 2008 17:18:35 +0000

42 lines

New Thread

HEA-NOC/20080224-1 [OPEN] Leased lines to two clients in the Mount Street area are down

HEA-NOC/20080224-1 [OPEN] Leased lines to two clients in the Mount Street area are down

Brian Boyle <[log in to unmask]>

Sun, 24 Feb 2008 09:06:11 +0000

33 lines

New Thread

HEA-NOC/20080224-1 [UPDATE] Leased lines to two clients in the Mount Street area are down

HEA-NOC/20080224-1 [UPDATE] Leased lines to two clients in the Mount Street area are down

Brian Boyle <[log in to unmask]>

Sun, 24 Feb 2008 12:20:43 +0000

37 lines

New Thread

HEA-NOC/20080225-1 [CLOSE] Setup mysql replication between primary and standby webserver

HEA-NOC/20080225-1 [CLOSE] Setup mysql replication between primary and standby webserver

Rob Gallagher <[log in to unmask]>

Mon, 25 Feb 2008 15:28:02 +0000

51 lines

New Thread

HEA-NOC/20080225-1 [OPEN] Setup mysql replication between primary and standby webserver

HEA-NOC/20080225-1 [OPEN] Setup mysql replication between primary and standby webserver

Rob Gallagher <[log in to unmask]>

Mon, 25 Feb 2008 09:45:53 +0000

45 lines

New Thread

HEA-NOC/20080225-10 [OPEN] IT Carlow Migration to New IP Equipment

HEA-NOC/20080225-10 [OPEN] IT Carlow Migration to New IP Equipment

Garwin Liu <[log in to unmask]>

Mon, 25 Feb 2008 14:32:27 +0000

44 lines

New Thread

HEA-NOC/20080225-10 [UPDATE] IT Carlow Migration to New IP Equipment

HEA-NOC/20080225-10 [UPDATE] IT Carlow Migration to New IP Equipment

Garwin Liu <[log in to unmask]>

Thu, 28 Feb 2008 08:20:41 +0000

54 lines

New Thread

HEA-NOC/20080225-12 [CLOSE] Start up listserv on second virtual machine

HEA-NOC/20080225-12 [CLOSE] Start up listserv on second virtual machine

Rob Gallagher <[log in to unmask]>

Mon, 25 Feb 2008 17:36:55 +0000

38 lines

New Thread

HEA-NOC/20080225-12 [OPEN] Start up listserv on second virtual machine

HEA-NOC/20080225-12 [OPEN] Start up listserv on second virtual machine

Rob Gallagher <[log in to unmask]>

Mon, 25 Feb 2008 15:50:42 +0000

34 lines

New Thread

HEA-NOC/20080225-8 [CLOSE] Commodity IP service maintenance

HEA-NOC/20080225-8 [CLOSE] Commodity IP service maintenance

Daniel Lete <[log in to unmask]>

Tue, 26 Feb 2008 09:52:54 +0000

39 lines

New Thread

HEA-NOC/20080225-8 [OPEN] Commodity IP service maintenance

HEA-NOC/20080225-8 [OPEN] Commodity IP service maintenance

Daniel Lete <[log in to unmask]>

Mon, 25 Feb 2008 14:19:25 +0000

33 lines

New Thread

HEA-NOC/20080225-9 [CLOSE] AIT Migration to New IP Equipment

HEA-NOC/20080225-9 [CLOSE] AIT Migration to New IP Equipment

Garwin Liu <[log in to unmask]>

Thu, 28 Feb 2008 08:38:23 +0000

52 lines

New Thread

HEA-NOC/20080225-9 [OPEN] AIT Migration to New IP Equipment

HEA-NOC/20080225-9 [OPEN] AIT Migration to New IP Equipment

Garwin Liu <[log in to unmask]>

Mon, 25 Feb 2008 14:30:46 +0000

43 lines

New Thread

HEA-NOC/20080225-9 [UPDATE] AIT Migration to New IP Equipment

HEA-NOC/20080225-9 [UPDATE] AIT Migration to New IP Equipment

Garwin Liu <[log in to unmask]>

Wed, 27 Feb 2008 09:01:55 +0000

49 lines

New Thread

HEA-NOC/20080227-3 [OPEN] BT Maintenance

HEA-NOC/20080227-3 [OPEN] BT Maintenance

Daniel Lete <[log in to unmask]>

Wed, 27 Feb 2008 09:22:27 +0000

34 lines

New Thread

HEA-NOC/20080227-4 [OPEN] BT Maintenance

HEA-NOC/20080227-4 [OPEN] BT Maintenance

Daniel Lete <[log in to unmask]>

Wed, 27 Feb 2008 09:22:34 +0000

31 lines

New Thread

HEA-NOC/20080227-6 [OPEN] GTI (Georgia Tech Institute) Migration to New IP equipment

HEA-NOC/20080227-6 [OPEN] GTI (Georgia Tech Institute) Migration to New IP equipment

Cormac O'Cianain <[log in to unmask]>

Wed, 27 Feb 2008 12:13:33 +0000

33 lines

New Thread

HEA-NOC/20080227-8 [CLOSE] Scheduled power work in TCD PoP

HEA-NOC/20080227-8 [CLOSE] Scheduled power work in TCD PoP

Louis Twomey <[log in to unmask]>

Fri, 29 Feb 2008 16:07:55 +0000

46 lines

New Thread

HEA-NOC/20080227-8 [OPEN] Scheduled power work in TCD PoP

HEA-NOC/20080227-8 [OPEN] Scheduled power work in TCD PoP

Louis Twomey <[log in to unmask]>

Wed, 27 Feb 2008 15:48:55 +0000

42 lines

New Thread

HEA-NOC/20080228-13 [CLOSE] Loss of Connectivity - HRB

HEA-NOC/20080228-13 [CLOSE] Loss of Connectivity - HRB

Garwin Liu <[log in to unmask]>

Fri, 29 Feb 2008 09:22:57 +0000

32 lines

New Thread

HEA-NOC/20080228-13 [OPEN] Loss of Connectivity - HRB

HEA-NOC/20080228-13 [OPEN] Loss of Connectivity - HRB

Garwin Liu <[log in to unmask]>

Thu, 28 Feb 2008 19:15:37 +0000

28 lines

New Thread

HEA-NOC/20080228-4 [CLOSE] HEAnet Office Phone System Maintenance

HEA-NOC/20080228-4 [CLOSE] HEAnet Office Phone System Maintenance

Brian Nisbet <[log in to unmask]>

Thu, 28 Feb 2008 13:09:21 +0000

37 lines

New Thread

HEA-NOC/20080228-4 [OPEN] HEAnet Office Phone System Maintenance

HEA-NOC/20080228-4 [OPEN] HEAnet Office Phone System Maintenance

Brian Nisbet <[log in to unmask]>

Thu, 28 Feb 2008 11:53:11 +0000

33 lines

New Thread

HEA-NOC/20080229-10 [OPEN] St Angela's link flapped

HEA-NOC/20080229-10 [OPEN] St Angela's link flapped

Rob Gallagher <[log in to unmask]>

Fri, 29 Feb 2008 20:37:24 +0000

30 lines

New Thread

HEA-NOC/20080229-4 [OPEN] Listserv services unavailable

HEA-NOC/20080229-4 [OPEN] Listserv services unavailable

Gareth Eason <[log in to unmask]>

Fri, 29 Feb 2008 11:41:07 +0000

27 lines

New Thread

HEA-NOC/20080229-4 [UPDATE] Listserv services unavailable

HEA-NOC/20080229-4 [UPDATE] Listserv services unavailable

Rob Gallagher <[log in to unmask]>

Fri, 29 Feb 2008 12:38:55 +0000

43 lines

HEA-NOC/20080229-4 [UPDATE] Listserv services unavailable

Gareth Eason <[log in to unmask]>

Fri, 29 Feb 2008 12:25:43 +0000

39 lines

New Thread

HEA-NOC/20080229-9 [OPEN] Cessation of ITnet STM-1 (Cork)

HEA-NOC/20080229-9 [OPEN] Cessation of ITnet STM-1 (Cork)

Owen Byrne <[log in to unmask]>

Fri, 29 Feb 2008 15:55:31 +0000

36 lines

New Thread

HEAnet Extended NOC Cover announcement...

HEAnet Extended NOC Cover announcement...

Gareth Eason <[log in to unmask]>

Thu, 7 Feb 2008 08:13:56 +0000

112 lines

New Thread

[OPEN] The servers hosting the video streaming service are down

[OPEN] The servers hosting the video streaming service are down

Gareth Eason <[log in to unmask]>

Tue, 19 Feb 2008 09:55:22 +0000

49 lines

New Thread

[UPDATE] Hardware failure of samhain, web services server

[UPDATE] Hardware failure of samhain, web services server

Gareth Eason <[log in to unmask]>

Tue, 19 Feb 2008 10:06:18 +0000

57 lines

Advanced Options


Options

Log In

Log In

Get Password

Get Password


Search Archives

Search Archives


Subscribe or Unsubscribe

Subscribe or Unsubscribe


Archives

March 2017
February 2017
January 2017
December 2016
November 2016
October 2016
September 2016
August 2016
July 2016
June 2016
May 2016
April 2016
February 2016
January 2016
December 2015
November 2015
October 2015
September 2015
August 2015
July 2015
June 2015
May 2015
April 2015
March 2015
February 2015
January 2015
December 2014
November 2014
October 2014
September 2014
August 2014
July 2014
June 2014
May 2014
April 2014
March 2014
February 2014
January 2014
December 2013
November 2013
October 2013
September 2013
August 2013
July 2013
June 2013
May 2013
April 2013
March 2013
February 2013
January 2013
December 2012
November 2012
October 2012
September 2012
August 2012
July 2012
June 2012
May 2012
April 2012
March 2012
February 2012
January 2012
December 2011
November 2011
October 2011
September 2011
August 2011
July 2011
June 2011
May 2011
April 2011
March 2011
February 2011
January 2011
December 2010
November 2010
October 2010
September 2010
August 2010
July 2010
June 2010
May 2010
April 2010
March 2010
February 2010
January 2010
December 2009
November 2009
October 2009
September 2009
August 2009
July 2009
June 2009
May 2009
April 2009
March 2009
February 2009
January 2009
December 2008
November 2008
October 2008
September 2008
August 2008
July 2008
June 2008
May 2008
April 2008
March 2008
February 2008
January 2008
December 2007
November 2007
October 2007
September 2007
August 2007
July 2007
June 2007
May 2007
April 2007
March 2007
February 2007
January 2007
December 2006
November 2006
October 2006
September 2006
August 2006
July 2006
June 2006
May 2006
April 2006
March 2006
February 2006
January 2006
December 2005
November 2005
October 2005
September 2005
August 2005
July 2005
June 2005
May 2005
April 2005
March 2005
February 2005
January 2005
December 2004
November 2004
October 2004
September 2004
August 2004
July 2004
June 2004
May 2004
April 2004
March 2004
February 2004
January 2004
December 2003
November 2003
October 2003
September 2003
August 2003
July 2003
June 2003
May 2003
April 2003
March 2003
February 2003
January 2003
December 2002
November 2002
October 2002
September 2002
August 2002
July 2002
June 2002
May 2002
April 2002
March 2002
February 2002
January 2002
December 2001
November 2001
October 2001
September 2001
August 2001
July 2001
June 2001
May 2001
April 2001
March 2001
February 2001
January 2001
December 2000
November 2000
October 2000
September 2000
August 2000
July 2000
June 2000
May 2000
April 2000
March 2000
February 2000
January 2000
December 1999
November 1999
October 1999
September 1999
August 1999
July 1999
June 1999
May 1999
April 1999
March 1999
February 1999
January 1999
December 1998
November 1998
October 1998
September 1998
August 1998
July 1998
June 1998
May 1998
April 1998
March 1998
February 1998
January 1998
December 1997
November 1997
October 1997
September 1997
August 1997
July 1997
June 1997
May 1997
April 1997
March 1997
February 1997
January 1997
December 1996
November 1996
October 1996
September 1996
August 1996
July 1996
June 1996
May 1996
April 1996
March 1996
February 1996
January 1996
December 1995
November 1995
October 1995
September 1995
August 1995
July 1995
June 1995
May 1995
April 1995
March 1995
February 1995
January 1995
December 1994
November 1994
October 1994
September 1994
August 1994
July 1994
June 1994
May 1994
April 1994
March 1994
February 1994
January 1994
December 1993
November 1993
October 1993
September 1993
August 1993
July 1993
June 1993

ATOM RSS1 RSS2



LISTSERV.HEANET.IE

CataList Email List Search Powered by the LISTSERV Email List Manager